# STRIDE: Sparse Techniques for Regression in Deep Gaussian Processes

### v1

@@ Abstract
Gaussian processes (GPs) have gained popularity as flexible machine learning models for regression and function approximation with an in-built method for uncertainty quantification. However, GPs suffer when the amount of training data is large or when the underlying function contains multi-scale features that are difficult to represent by a stationary kernel. To address the former, training of GPs with large-scale data is often performed through inducing point approximations (also known as sparse GP regression (GPR)), where the size of the covariance matrices in GPR is reduced considerably through a greedy search on the data set. To aid the latter, deep GPs have gained traction as hierarchical models that resolve multi-scale features by combining multiple GPs. Posterior inference in deep GPs requires a sampling or, more usual, a variational approximation. Variational approximations lead to large-scale stochastic, non-convex optimisation problems and the resulting approximation tends to represent uncertainty incorrectly. In this work, we combine variational learning with MCMC to develop a particle-based expectation-maximisation method to simultaneously find inducing points within the large-scale data (variationally) and accurately train the GPs (sampling-based). The result is a highly efficient and accurate methodology for deep GP training on large-scale data. We test our method on standard benchmark problems.
@@ 1 Introduction
Gaussian process regression is a popular Bayesian methodology in supervised learning, non-parametric statistics, and function approximation, with a variety of favourable features such as tractable inference and uncertainty quantification. When GPs are not sufficiently flexible, we can construct deep GPs by, e.g., sequentially composing a number of otherwise independent GPs. Deep GPR retains many features of GPR, but importantly makes posterior inference much more computationally challenging. Sampling-based techniques for deep GPR require repeated GPR steps – infeasible if the number of observed data sets is large. In the present work, we propose STRIDE a method that combines Markov chain Monte Carlo (MCMC) with sparse GPR for an accurate and computationally efficient deep GPR. Indeed, we
@@
    discuss approximate sampling techniques on lower deep Gaussian process layers to reduce the overall complexity to linear in the size of the training dataset,
@@
After a literature review, we collect the technical background in conventional, deep, and sparse GPR in Section 2, before we derive STRIDE in Section 3, summarise our experiments and their results in Section 4, and conclude in Section 5.
Literature review.
Sparse GPR is a popular tool for solving regression problems involving large numbers of data points, see e.g. the works [41, 35, 7, 38] as well as the unifying perspectives [3, 30]. In STRIDE, we focus on the objective function described in [38] to select optimal inducing points.
In the context of deep GPR, the posterior needs to be approximated with samples, e.g. through MCMC [14, 24, 34, 23], or a variational approximation [10, 19, 9, 8]. Variational approximations are popular due to their scalability, but challenges remain as they require non-convex large-scale optimisation, tend to misrepresent uncertainty, and are not yet fully understood theoretically [2]. In this work, we therefore focus on provably convergent MCMC methods for reliable uncertainty quantification. Sparse GPR techniques based on [38] have successfully been employed in variational approximations of deep GPs [10, 11], as well as the related setting of hierarchical GPs with hyper-parameters [32].
[…]
Convolution
[…]
Composition
Deep GPs can be constructed in a variety of ways, typically based on a hierarchical approach to non-stationary GP modelling [14, 33]. When combining the individual GP layers in said hierarchical fashion, the most prominent example is the composition architecture as described in the seminal papers [10, 15]. The convolution architecture, on the other hand, uses the individual GP layers to define a non-stationary correlation length for the next layer [31, 14]. We illustrate both architectures in Figure 1. A synthesis of composition and convolution architectures was recently proposed in [12].
STRIDE is based on the expectation-maximisation (EM) algorithm [13, 36]. We also refer to [25] for variants of EM, including Monte Carlo EM [39] which is used in this work. While we combine preconditioned Crank–Nicolson (pCN) MCMC [6] for the expectation step and a greedy method for the maximisation step, previous works have combined Langevin sampling and gradient flows [20, 22] in a similar context.
@@ 2 Gaussian process regression and extensions
In the following we consider the following supervised learning problem: we are given data pairs $(x_{i},y_{i})_{i=1}^{n}\in(\mathbb{R}^{d}\times\mathbb{R})^{n}$ through $f^{\dagger}(x_{i})+\varepsilon_{i}=y_{i}$, where $f^{\dagger}$ is an unknown function and $\varepsilon_{1},...,\varepsilon_{n}\sim\mathrm{N}(0,\gamma^{2})$ i.i.d. is observational noise. Our goal is to recover the function $f^{\dagger}$ from the observational data. GPR is a natural Bayesian strategy to approach problems of this form. Given a prior $\mathrm{GP}(0,k(\cdot,\cdot))$, we obtain the posterior $\mathrm{GP}(\mu^{y}(\cdot),k^{y}(\cdot,\cdot))$, with
$\mu^{y}(\cdot)$ $=k(\cdot,\mathbf{x}_{n})\mathopen{}\mathclose{{}\left(k(\mathbf{x}_{n},\mathbf{x}_{n})+\gamma^{2}\mathrm{Id}}\right)^{-1}\mathbf{y}_{n}\>,$ (1)
$k^{y}(\cdot,\cdot)$ $=k(\cdot,\cdot)-k(\cdot,\mathbf{x}_{n})\mathopen{}\mathclose{{}\left(k(\mathbf{x}_{n},\mathbf{x}_{n})+\gamma^{2}\mathrm{Id}_{n}}\right)^{-1}k(\mathbf{x}_{n},\cdot)\>.$
Here, the bold letters refer to the full data sets $\mathbf{x}_{n}:=(x_{i})_{i=1}^{n}$ and $\mathbf{y}_{n}:=(y_{i})_{i=1}^{n}$. When kernels are applied to these sets, we obtain kernel matrices/vectors.
In the following, we discuss two extensions of GPR: deep GPR in cases where GPs are not sufficiently flexible (Subsections 2.1, 2.2) and sparse GPR as a computationally efficient approximation of GPR.(Subsection 2.3).
@@ 2.1 Deep Gaussian processes
A sequence of random functions $(f_{\ell})_{0\leq\ell\leq\ L}$ is called a deep Gaussian process with $L+1$ layers if $f_{0}$ is a Gaussian process, and if $f_{\ell}$ is conditionally Gaussian for $\ell\geq 1$, i.e.
$\pi_{f_{\ell}|f_{\ell-1}=u_{\ell-1}}=\mathrm{GP}(0,k_{\ell}(\cdot,\cdot\,|u_{\ell-1})),$
@@
The properties of a deep GP strongly depend on the choice of the conditional covariance kernel $k_{\ell}$, which determines its architecture. We consider two approaches here, namely, the convolution architecture described in [27, 14], as well as the composition/warping architecture [10, 14, 34].
@@
Covariance kernels are often constructed in the following way: we define an isotropic correlation function $\rho:\mathbb{R}\rightarrow\mathbb{R}$ and then obtain a covariance kernel through $k(x,x^{\prime})=\sigma^{2}\rho(\|x-x^{\prime}\|\>/\>\lambda)$ for some marginal variance $\sigma^{2}>0$ and correlation length $\lambda>0$. The correlation length determines the strength of long-range correlation in a GP and, thus, the oscillatory behaviour of its samples. In the convolution architecture, we now use a spatially-dependent value for $\lambda$ that is given through the realisation $u_{\ell-1}$ of the next-lower deep GP layer $f_{\ell-1}$. A spatially variable correlation length allows us to describe multi-scale behaviour in a function, e.g., in images. We usually set $\lambda(x)\coloneqq h(u_{\ell-1}(x))$, where $h:\>\mathbb{R}\rightarrow\mathbb{R}_{+}$ is some non-negative transformation of the hidden layer values $u_{\ell-1}$. More precisely, in a multi-layer setting, we give the covariance function $k_{\ell}$ as,
$k_{\ell}^{\mathrm{conv}}(x,x^{\prime}\,|\,u_{\ell-1})$ $=\sigma^{2}\,\phi(x,x^{\prime})\,\rho\mathopen{}\mathclose{{}\left(\frac{2\,\|x-x^{\prime}\|}{h(u_{\ell-1}(x))+h(u_{\ell-1}(x^{\prime}))}}\right),$
@@
$\phi(x,x^{\prime})=\frac{2^{d/2}\,h(u_{\ell-1}(x))^{d/4}\,h(u_{\ell-1}(x^{\prime}))^{d/4}}{(h(u_{\ell-1}(x))+h(u_{\ell-1}(x^{\prime})))^{d/2}},\qquad h(z)=\min(\max(z,u^{\min}),u^{\mathrm{max}})^{2}\>.$
In the above, the pre-factor $\phi(x,x^{\prime})$ ensures that the resulting covariance function remains positive semi-definite. The scalar transformation $h:\mathbb{R}\rightarrow\mathbb{R}^{+}$ given here is an example that worked well in our experiments, with $u^{\min},u^{\mathrm{max}}$ as tunable parameters. Other functions that result in positive correlation length values would also be valid. Note that the correlation length can in principle be chosen to be matrix-valued, but we restrict ourselves to scalars in this work, see [27, 14].
@@
$k_{\ell}^{\mathrm{comp}}(x,\,x^{\prime}\,|\,u_{\ell-1})$ $=k(u_{\ell-1}(x),\,u_{\ell-1}(x^{\prime}))\>.$
In this case, some of the layers may be multi-output GPs.

### v2

# Sparse Techniques for Regression in Deep Gaussian Processes

@@ Abstract
Gaussian processes (GPs) have gained popularity as flexible machine learning models for regression and function approximation with an in-built method for uncertainty quantification. However, GPs suffer when the amount of training data is large or when the underlying function contains multi-scale features that are difficult to represent by a stationary kernel. To address the former, training of GPs with large-scale data is often performed through inducing point approximations, also known as sparse GP regression (GPR), where the size of the covariance matrices in GPR is reduced considerably through a greedy search on the data set. To aid the latter, deep GPs have gained traction as hierarchical models that resolve multi-scale features by combining multiple GPs. Posterior inference in deep GPs requires a sampling or, more usual, a variational approximation. Variational approximations lead to large-scale stochastic, non-convex optimisation problems and the resulting approximation tends to represent uncertainty incorrectly. In this work, we combine variational learning with MCMC to develop a particle-based expectation-maximisation method to simultaneously find inducing points within the large-scale data (variationally) and accurately train the deep GPs (sampling-based). The result is a highly efficient and accurate methodology for deep GP training on large-scale data. We test our method on standard benchmark problems.
@@ 1 Introduction
Gaussian process regression (GPR) is a popular Bayesian methodology in supervised learning, non-parametric statistics, and function approximation, with a variety of favourable features such as tractable inference and uncertainty quantification. The GP priors used in GPR are often parametrised by a stationary covariance kernel, which may fail to accurately represent non-stationary features of the quantity to be approximated. A possible remedy is to employ a non-stationary deep GP prior instead. Such deep GPs can be constructed by, e.g., sequentially composing a number of otherwise independent GPs. Deep GPR retains many features of GPR, but importantly makes posterior inference much more computationally challenging. Sampling-based techniques for deep GPR require repeated GPR steps, which is infeasible if the number of observed data points is large. As a result, an efficient strategy for performing these GPR steps is needed to obtain a viable deep GPR method. In stationary GPR, sparse GPR methods have seen great success in reducing the computational complexity, particularly for large datasets. Sparse GPR relies on a low-rank approximation of the exact GP covariance. Typically, this low-rank approximation depends on a set of inducing points, that is, a set of points that is much smaller than the original dataset, but can be regarded as representative of the entire dataset.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Example of reconstructing a one-dimensional function using GP regression, sparse GP regression, and STRIDE, respectively.
We illustrate the need for a methodology that combines deep and sparse GP regression techniques with a small example. We aim to reconstruct a one-dimensional ground truth function, shown in the upper left plot in Figure˜1, observed at 200 locations distributed in a uniform grid on $[0,1]$. The observations are furthermore polluted with Gaussian noise with a small variance of $0.02^{2}$. This ground truth is challenging in the context of GPR for the fact that it is constructed from a very smooth part on the left, and a part that contains a series of discontinuous jumps to the right. That is, we expect a stationary covariance kernel to struggle to represent this ground truth function. We show the posterior mean obtained from standard GP regression in the upper right plot in Figure˜1, using optimised hyper-parameters and a Gaussian covariance kernel. It is clear that we recover the left, smooth side of the ground truth very accurately, while the reconstructed function exhibits over-shooting at the jumps on the right. However, in applications involving large-scale datasets, using standard GPR is typically infeasible. Instead, a popular choice is the use of sparse GPR methods, such as described in [titsias2009variational, titsias2009report]. We employ a greedy algorithm to choose the inducing points, and optimise the hyper-parameters each time a new point is selected. We present the result of this approach for a number of 20 inducing points in the lower left plot, with the inducing point locations marked as black crosses. In this example, it becomes clear that the computational speed-up we gain from using sparse GPR comes at the cost of reducing the accuracy of the reconstruction. Namely, the jumps on the right-hand side of the ground truth function are recovered rather poorly by the sparse GPR mean. The goal of the method introduced in this paper is to improve the accuracy of sparse GPR, while maintaining its scalability properties. In our small example, we show the reconstruction obtained with STRIDE using a 3-layer deep GP architecture in the lower right plot in Figure˜1. Here, we can observe that, with the same number of inducing points, we are able to represent the ground truth more accurately. Namely, the slopes of the reconstruction around the jump locations are larger and thus more consistent with the ground truth, while slightly reducing the overshoot occurring at the same locations compared to standard and sparse GPR. We provide more details on this one-dimensional example in Section˜4.1.
Sparse GPR is a popular tool for solving regression problems involving large numbers of data points, see e.g. the works [williams2000using, smola2000sparse, csato2002sparse, titsias2009variational] as well as the unifying perspectives [bui2017unifying, quinonero2005unifying]. In STRIDE, we focus on the objective function described in [titsias2009variational] to select optimal inducing points. This is only one of many possible approaches to choosing inducing points. In [bui2017unifying], a new framework is introduced that generalises the free variational energy used in [titsias2009variational] as well as the deterministic, fully independent, and partially independent training conditionals summarised in [quinonero2005unifying]. A good overview and introduction to this and other popular inducing point selection strategies in sparse GPR is given in [bui2017unifying, quinonero2005unifying]. As sparse GPR methods typically scale as $\mathcal{O}(nm^{3})$ in computational complexity (and as $\mathcal{O}(nm)$ in memory), where $n$ is the number of data points and $m$ the number of inducing points, [wilson2015kernel] introduce a method relying on Kronecker and Toeplitz algebra to improve the scaling w.r.t. $m$ in order to obtain an efficient method when large values of $m$ are needed to retain accuracy.
In the context of deep GPR, the posterior needs to be approximated with samples, e.g. through MCMC [dunlop18, monterrubio2020posterior, sauer2023active, latz2025deep], or a variational approximation [damianou2013deep, hensman2014nested, dai2015variational, cutajar2017random]. Variational approximations are popular due to their scalability, but challenges remain as they require non-convex large-scale optimisation, tend to misrepresent uncertainty [jankowiak2020parametric], and are not yet fully understood theoretically [Blei03042017]. In this work, we therefore focus on provably convergent MCMC methods for reliable uncertainty quantification. Sparse GPR techniques based on [titsias2009variational] have successfully been employed in variational approximations of deep GPs [damianou2013deep, damianou2011variational]. On the other hand, [rossi2021sparse] introduced a framework for inferring inducing points together with hyper-parameters in a Bayesian fashion, including an extension to deep GPs. The method we present in this work can be seen as taking a middle route: while we employ an MCMC algorithm to obtain samples of the hidden layers, we choose inducing points through an optimisation approach, resulting in an expectation-maximisation (EM) algorithm [dempster1977maximum, sundberg1974maximum, Ng2012, wei1990monte]. This methodology can also be linked to the work of [johnston2025interacting, kuntz23a] who combined Langevin sampling and gradient flows in a similar context. Other works go beyond more classical GP tools by using a stochastic differential equation to sample inducing points [xu2024sparse], or a deep neural network to model non-stationarity in a GP [wilson2016deep].
[…]
Convolution
[…]
Composition
Deep GPs can be constructed in a variety of ways, typically based on a hierarchical approach to non-stationary GP modelling [dunlop18, sauer2023non]. When combining the individual GP layers in said hierarchical fashion, the most prominent example is the composition architecture as described in the seminal papers [damianou2013deep, duvenaud2014avoiding]. The convolution architecture, on the other hand, uses the individual GP layers to define a non-stationary correlation length for the next layer [roininen2019hyper, dunlop18]. We illustrate both architectures in Figure˜2. A synthesis of composition and convolution architectures was recently proposed in [de2023thin]. Another notable way of constructing deep GPs is the use of non-stationary covariance operators that are defined via the realisations of the next-lower layers [latz2025deep, dunlop18].
In the present work, we propose STRIDE (short for Sparse Techniques for Regression in Deep Gaussian Processes), a method that combines Markov chain Monte Carlo (MCMC) with sparse GPR for an accurate and computationally efficient deep GPR. Indeed, we
@@
    discuss approximate sampling techniques on lower deep Gaussian process layers to reduce the overall complexity to be linear in the size of the training dataset,
@@
Structure.
Following this introduction, we collect the technical background on conventional, deep, and sparse GPR in Section˜2. Based on this, we combine deep and sparse GP techniques to derive STRIDE in Section˜3. We describe the behaviour and performance of our algorithm in a range of numerical experiments, presenting their setup and results in Section˜4. Finally, we conclude in Section˜5.
@@ 2 Gaussian process regression and extensions
In the following, we consider a supervised learning problem: we are given data pairs $(\mathbf{x}_{i},y_{i})_{i=1}^{n}\in(\mathbb{R}^{d}\times\mathbb{R})^{n}$ through $f^{\dagger}(\mathbf{x}_{i})+\varepsilon_{i}=y_{i}$, where $f^{\dagger}$ is an unknown function and $\varepsilon_{1},...,\varepsilon_{n}\sim\mathrm{N}(0,\gamma^{2})$ is observational noise assumed to be independent and identically distributed. We denote the full dataset divided into observation locations and values $X:=(\mathbf{x}_{i})_{i=1}^{n}$ and $\mathbf{y}:=(y_{i})_{i=1}^{n}$. Our goal is to recover the function $f^{\dagger}$ from the observational data. GPR is a natural Bayesian strategy to approach problems of this form. Given a prior $\mathcal{GP}(0,k)$, we obtain the posterior $\mathcal{GP}(\mu^{\mathbf{y}},k^{\mathbf{y}})$, with
$\mu^{\mathbf{y}}$ $=k(\cdot,X)\mathopen{}\mathclose{{\left(k(X,X)+\gamma^{2}\mathrm{Id}_{n}}}\right)^{-1}\mathbf{y}\>,$ (1)
$k^{\mathbf{y}}$ $=k-k(\cdot,X)\mathopen{}\mathclose{{\left(k(X,X)+\gamma^{2}\mathrm{Id}_{n}}}\right)^{-1}k(X,\cdot)\>.$
In the above, we slightly abuse notation by using $X$ as an argument of the GP covariance kernel $k$. We use $k(X,X)$ to denote the matrix resulting from evaluating $k(\mathbf{x}_{i},\mathbf{x}_{j})$ for each pair of indices $1\leq i,j\leq n$. The expressions $k(\cdot,X)$ and $k(X,\cdot)$ are maps from $\mathbb{R}^{d}$ into $\mathbb{R}^{n}$ and contain entries $k(\cdot,\mathbf{x}_{i})$ for $1\leq i\leq n$, given as a row vector and a column vector, respectively.
In the following, we discuss two extensions of GPR: deep GPR in cases where GPs are not sufficiently flexible (Subsections 2.1, 2.2) and sparse GPR as a computationally efficient approximation of GPR (Subsection 2.3).
@@ 2.1 Deep Gaussian processes
A sequence of random functions $(f_{\ell})_{0\leq\ell\leq\ L}$ is called a deep Gaussian process with $L+1$ layers if $f_{0}$ is a Gaussian process, and if $f_{\ell}$ is conditionally Gaussian for $1\leq\ell\leq L$, i.e.
$\pi_{f_{\ell}|f_{\ell-1}=u_{\ell-1}}=\mathcal{GP}(0,k_{\ell}(\cdot,\cdot\,|\,u_{\ell-1}))\>,$
@@
The properties of a deep GP strongly depend on the choice of the conditional covariance kernel $k_{\ell}$, which determines its architecture. We consider two approaches here, namely, the convolution architecture described in [paciorek03, dunlop18], as well as the composition/warping architecture [damianou2013deep, dunlop18, sauer2023active]. We illustrate both constructions in Figure˜2, which we refer to in what follows. In this figure, we show the estimated posterior mean for each layer, conditioned on data from a piece-wise constant ground truth function.
@@
Covariance kernels are often constructed in the following way: we define an isotropic correlation function $\rho:\mathbb{R}\rightarrow\mathbb{R}$ and then obtain a covariance kernel through $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\rho(\|\mathbf{x}-\mathbf{x}^{\prime}\|\>/\>\lambda)$ for some marginal variance $\sigma^{2}>0$ and correlation length $\lambda>0$. The correlation length determines the strength of long-range correlation in a GP and, thus, the oscillatory behaviour of its samples. In the convolution architecture, we now use a spatially-dependent value for $\lambda$ that is given through the realisation $u_{\ell-1}$ of the next-lower deep GP layer $f_{\ell-1}$. A spatially variable correlation length allows us to describe multi-scale behaviour in a function, e.g., in images. We usually set $\lambda(\mathbf{x})\coloneqq h(u_{\ell-1}(\mathbf{x}))$, where $h:\>\mathbb{R}\rightarrow\mathbb{R}_{+}$ is some non-negative transformation of the hidden layer values $u_{\ell-1}$. More precisely, in a multi-layer setting, we give the covariance function $k_{\ell}$ as,
$k_{\ell}^{\mathrm{conv}}$ $(\mathbf{x},\mathbf{x}^{\prime}\,|\,u_{\ell-1})=\sigma^{2}\,\phi(\mathbf{x},\mathbf{x}^{\prime})\,\rho\mathopen{}\mathclose{{\left(\frac{2\,\|\mathbf{x}-\mathbf{x}^{\prime}\|}{h(u_{\ell-1}(\mathbf{x}))+h(u_{\ell-1}(\mathbf{x}^{\prime}))}}}\right),$
@@
$\phi(\mathbf{x},\mathbf{x}^{\prime})=\frac{2^{d/2}\,h(u_{\ell-1}(\mathbf{x}))^{d/4}\,h(u_{\ell-1}(\mathbf{x}^{\prime}))^{d/4}}{(h(u_{\ell-1}(\mathbf{x}))+h(u_{\ell-1}(\mathbf{x}^{\prime})))^{d/2}},\qquad h(z)=\min(\max(z,u^{\mathrm{min}}),u^{\mathrm{max}})^{2}\>.$
In the above, the pre-factor $\phi(\mathbf{x},\mathbf{x}^{\prime})$ ensures that the resulting covariance function remains positive semi-definite [paciorek03]. The scalar transformation $h:\mathbb{R}\rightarrow\mathbb{R}^{+}$ given here is an example that worked well in our experiments, with $u^{\mathrm{min}},u^{\mathrm{max}}$ as tunable parameters. Other functions that result in positive correlation length values would also be valid. Note that the correlation length can in principle be chosen to be matrix-valued, but we restrict ourselves to scalars in this work, see [paciorek03, dunlop18]. In Figure˜2, this construction has the effect that we observe small mean values of the hidden layers close to the jump locations of the ground truth. Due to our choice of $h$, the small hidden layer values translate into a small local length scale in these areas.
@@
$k_{\ell}^{\mathrm{comp}}(\mathbf{x},\,\mathbf{x}^{\prime}\,|\,u_{\ell-1})$ $=k(u_{\ell-1}(\mathbf{x}),\,u_{\ell-1}(\mathbf{x}^{\prime}))\>.$
In this case, some of the layers may be multi-output GPs. This is also the case in our example in Figure˜2, where we choose GPs with two-dimensional output as the hidden layers. The output values of one hidden layer serve as the spatial input of the next. The local length scale is encoded in the slope of the next-lower layer. In this example, this means that we obtain large slopes in the hidden layer mean values in areas where we require a small length scale to accurately recover the jumps of the ground truth function.
One drawback of the composition construction is that it can suffer from mode collapse. That is, when a layer $u_{\ell-1}$ is not injective, distinct spatial locations $\mathbf{x}$, $\mathbf{x}^{\prime}$ can be mapped to the same value $u_{\ell-1}(\mathbf{x})=u_{\ell-1}(\mathbf{x}^{\prime})$, meaning that these points are fully correlated in the remainder of the deep GP layers. This issue is quite prominent in one dimension, and has also been described theoretically [duvenaud2014avoiding]. As a way to avoid mode collapse in this setting, we adapt the composition construction to enforce strictly monotonous layer inputs. We define this injective non-stationary kernel as,
$k_{\ell}^{\mathrm{inj}}(x,x^{\prime}\,|\,u_{\ell-1})$ $=k(g(u_{\ell-1};x),g(u_{\ell-1};x^{\prime}))$ (2)
for $x\in[0,1]$, where
$g(u_{\ell-1};x)$ $=\int_{0}^{x}\max\mathopen{}\mathclose{{\left(u_{\ell-1}(z),u^{\mathrm{min}}}}\right)^{2}\,\mathrm{d}z\>.$
See also [barnett2025monotonic] for a similar construction. Clearly, this setup only works in one dimension. Through the choice of $u^{\mathrm{min}}>0$, we can control the maximum correlation length that can occur in the deep GP construction. Effectively, it ensures that $g(u_{\ell-1};x)\geq(u^{\mathrm{min}})^{2}\,x$, which puts a constraint on the minimum positive slope of the warping $g$.